OpenStudy (anonymous):

What is the solution to the equation? t(7 – 2) = 9 • 7 – 9 • 2

OpenStudy (anonymous):

t=9/5 or t=1.8

OpenStudy (anonymous):

@Liamnomnom, read the terms of service, please.

OpenStudy (anonymous):

@bloopman No

OpenStudy (anonymous):

@Liamnomnom yes.

OpenStudy (♪chibiterasu):

It's the Code of Conduct that states you cannot blurt out the answer, not the Terms of Service.

OpenStudy (anonymous):

@bloopman Blocked :)

OpenStudy (♪chibiterasu):

You blocked him because he told you to read the rules? >_>

OpenStudy (anonymous):

@♪Chibiterasu I helped someone who asked for help, shoot me.

OpenStudy (whpalmer4):

\[t(7-2) = 9*7-9*2\]If you factor the right hand side, you get \[t(7-2) = 9(7-2)\] What does that suggest to you?

OpenStudy (♪chibiterasu):

Help is when you guide them through the question, here you are helping them cheat and you are not making it any better.

OpenStudy (♪chibiterasu):

whpalmer is actually helping, you see?

OpenStudy (anonymous):

@♪Chibiterasu can you report me? I tried deleting my account but I can't figure out how.

OpenStudy (anonymous):

I already have the answer so..

OpenStudy (whpalmer4):

So close the question, then...

OpenStudy (anonymous):

I know what to do. you dont have to tell me..

OpenStudy (♪chibiterasu):

This sas tho.

OpenStudy (anonymous):

The ToS also has a few words about conduct.

OpenStudy (♪chibiterasu):

But the Code of Conduct has more. :P